OpenStudy (anonymous):

Given the function f(x) = 3(x+2) - 4, solve for the inverse function when x = 2. I need help on this question!! I get really confused when it comes to the inverse of an equation.

OpenStudy (loser66):

solve 3(x+2)-4 =2, what do you get ? x =??

OpenStudy (anonymous):

I keep getting 3x=0. Am I supposed to use the distributive property first?

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

3(x +2) -4 =2 3(x+2) =6 x+2 = 2 x =0, get this part??

OpenStudy (anonymous):

That makes more sense.

OpenStudy (loser66):

Now, approach the problem on this way: \(f(\color{red}{x}) = a\\f^{-}(a) =\color{red}{x}\) Ok then??

OpenStudy (loser66):

You just tell me whether you got the previous post or not. Your problem is not done yet.

OpenStudy (anonymous):

I still don't really get it.

OpenStudy (anonymous):

I get that you flip the two. So if x is zero what would the a symbolize?

OpenStudy (loser66):



OpenStudy (loser66):